OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 3x + 5y = 3 x + 2y = 0 A. (1, 1) B. (6, –3) C. (6, 1) D. (1, –3)

OpenStudy (e.cociuba):

First solve for x in the 2nd equation by subtracting 2y on both sides of the equation x + 2y = 0 x = -2y

OpenStudy (e.cociuba):

Now substitute what x equals in the 1st equation 3x + 5y = 3 3(-2y) + 5y = 3 -6y + 5y = 3 -y = 3 y = -3

OpenStudy (anonymous):

B.?

OpenStudy (e.cociuba):

Correct :)

OpenStudy (primeralph):

No.

OpenStudy (anonymous):

Thank you!

OpenStudy (primeralph):



OpenStudy (e.cociuba):

Lol too funny :P No problemo! :)